Jumper cables used to start a stalled vehicle often carry a 65 A current. How strong is the magnetic field 17 cm from one cable?
______T

Compare to the earth's magnetic field (5.5 ✕ 10-5 T).
______%
Answer #1

B=u*I/2*pi*r

I=65 A

r=0.17 m

putting the values we got

B=7.65*10^(-5) T

COmpared to Earth

Percentage=(7.65/5.5)*100 =139.1 %